OpenStudy (anonymous):

Challenge: The distance around the middle of a sphere (the circumference) equals 37.699 in. What is the surface area of the sphere?

OpenStudy (anonymous):

I got 4,464.88 in^2.

OpenStudy (anonymous):

Is it correct?

OpenStudy (anonymous):

@dan815 @pooja195 @ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (jdoe0001):

yeap \(\bf \textit{surface of a sphere}=4\pi r^2\qquad r=\cfrac{37.699}{2}\qquad 4\cdot \pi \cdot \left( \cfrac{37.699}{2} \right)^2 \approx 4464.88\)

OpenStudy (jdoe0001):

\(\bf \textit{surface of a sphere}=4\pi r^2\qquad r=\cfrac{37.699}{2}\qquad \\ \quad \\ 4\cdot \pi \cdot \left( \cfrac{37.699}{2} \right)^2 \approx 4464.88\)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

What we did earlier was wrong

OpenStudy (anonymous):

C = 2πr 37.699 = 2πr 37.699 / 2π = r 59.22 ≈ r The radius is approximately 59.22 inches. SA = 4π59.22^2 SA ≈ 44070.37 The surface area is approximately 44,070.37 in^2.

OpenStudy (anonymous):

That's my work for my revision

OpenStudy (loser66):

First off, you need find r Yes

OpenStudy (loser66):

correct

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

what is correct?? XD

OpenStudy (loser66):

The last one.

OpenStudy (anonymous):

heey, @jdoe0001 , what we did earlier was wrong, we thought the circumference was the diameter and worked from there

OpenStudy (loser66):

while I tried to guide you, you posted the correct one, so that I just say yes.

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (jdoe0001):

hmmm the distance around the middle of a sphere, sounds to me like the diameter :)

OpenStudy (loser66):

@jdoe0001 Hey, invisible friend, no more invisible??? hahaha

OpenStudy (anonymous):

But in the question it says it's the circumference

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

It's the distance AROUND, not the distance from one end to another

OpenStudy (jdoe0001):

lemme be back in a bit

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u got it??

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

@sewandowski ... hold the mayo

OpenStudy (jdoe0001):

so the circumference is 37.699 we also know that the circumference of any circle is \(2\pi r\) thus we could say that \(\bf circumference = 37.699\qquad circumference=2\pi r\qquad thus \\ \quad \\ 37.699=2\pi r\implies {\color{brown}{ \cfrac{37.699}{2\pi } }}=r \\ \quad \\ \quad \\ \textit{surface area}=4\pi {\color{brown}{ r}}^2\to 4\pi \cdot \left( {\color{brown}{ \cfrac{37.699}{2\pi } }} \right)^2\to 4\pi \cfrac{37.699^2}{2^2\pi^2}\to \cancel{4\pi} \cfrac{37.699^2}{\cancel{4\pi}\pi }\)

OpenStudy (anonymous):

Heey, just finished the test, it was another answer :/

OpenStudy (anonymous):

Nevermind, thanks for the help but I need help in another problem, so I'll close this one and make another one. Thanks anyways!

OpenStudy (jdoe0001):

k

OpenStudy (mathstudent55):

\(r = \dfrac{37.699 ~in.}{2 \pi} = 5.99998 ~in.\) \(A = 4 \pi r^2 = 4 \times \pi \times (5.99998~in.)^2 = 452.39~in.^2\)

OpenStudy (mathstudent55):

@sewandowski Before you go look at this. You wrote above: C = 2πr 37.699 = 2πr 37.699 / 2π = r 59.22 ≈ r When you solve a problem, you need to do the operations, but you also need to confirm that answers make sense. If the circumference of a circle is approx. 38 in., how can the radius of the circle be 59.22 in.? It's impossible for the radius of a circle to be larger than the circumference.

OpenStudy (mathstudent55):

Your mistake above is this: How do you calculate \(\dfrac{37.699}{2 \pi} \) ? You can use your calculator and calculat4e first 2 * pi = 6.28 Then do 37.699/6.28 Or you can do 37.699 / 2 / pi You need to divide 37.699 by 2 and divide the result by pi, not multiply by pi like you did.